OpenStudy (anonymous):

Help pleasee will give medal and fan Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6. y = one thirdx + 14 thirds y = 3x + 10 y = -3x - 2 y = -one thirdx + 10 thirds

OpenStudy (anonymous):

helppp pllleeaassee

OpenStudy (anonymous):

slope intercept form is y=mx+b where m is the slope and b is the y intercept. Since you are looking for a parallel line you need an equation with the same slope as the first equation but a different y intercept So you plug your point into the given equation and change the y intercept from 6 to b y=-3x+6 4=-3(-2)+b 4=6+b 4-6=b -2=b So your new parallel equation is y=-3x-2

OpenStudy (anonymous):

wow thank you alot